The annual fixed cost of an equipment is $160,000, its generating income is
$1,200 per day & its operating cost per day is $240 for 12-hours per day and
300-days maximum per year. At 10% interest & 10years useful life evaluate this
equipment and find its current value.
1
Expert's answer
2020-05-26T10:52:30-0400

Taking into account the specified income and expenses for the year, the profit from the use of equipment will be


"(1200-240)\times 300=288000"

10%=0.1



"0.1\times288000=28800"


"28800\times 10-160000=128000"